# Lattice point enumeration of polytopes associated to integer compositions

### v2

@@ 3. Ehrhart and Zeta polynomials
This section offers two proofs of Theorem 1.2 and deduces a combinatorial interpretation of the $h^{\ast}$-polynomial of the composition polytope ${\mathcal{Q}}_{\sigma}$. As in the previous section, $\sigma=(r_{1},r_{2},\dots,r_{k})$ is a composition of $n$ and $s_{i}=r_{1}+r_{2}+\cdots+r_{i}$ for $0\leq i\leq k$, where $s_{0}:=0$.
@@
For any $m\in{\mathbb{N}}$, ${\rm Ehr}(Q_{\sigma},m)$ is equal to the number of points $(a_{1},a_{2},\dots,a_{n})\in{\mathbb{N}}^{n}$ such that
@@
Part of our second proof of Theorem 1.3 follows the proof of [12, Equation (33)], which is a formula for the number of lattice points of a Pitman–Stanley polytope. To avoid making too long of a digression, we assume familiarity with Sections 3 and 4 of [12] and only sketch this part of the proof.
@@
Let us consider the poset $\Omega_{\sigma}$ which consists of two disjoint chains $\alpha_{1}\prec\alpha_{2}\prec\cdots\prec\alpha_{n}$ and $\beta_{1}\prec\beta_{2}\prec\cdots\prec\beta_{n}$, together with the additional cover relations $\alpha_{s_{i}}\prec\beta_{s_{i}}$ for every $i\in[k]$. We let $C=\{\beta_{1},\beta_{2},\dots,\beta_{n}\}$ and ${\mathbf{u}}=(u_{1},u_{2},\dots,u_{n})$, where $u_{j}=jm$ for every $j\in[n]$ and where $m\in{\mathbb{N}}$ is fixed. In this situation, Pitman and Stanley [12, Section 3] have defined the order cone section ${\mathcal{C}}_{C}(\Omega_{\sigma},{\mathbf{u}})$ as the polytope of all order preserving maps $f:\Omega_{\sigma}\rightarrow{\mathbb{R}}_{\geq 0}$ such that $f(\beta_{j})=u_{j}$ for every $j\in[n]$. Setting $f_{j}=f(\alpha_{j})$ for every $j\in[n]$ and $f_{0}=0$ we have
@@
for every $i\in[k]$. The transformation $x_{j}=f_{j}-f_{j-1}$ for $j\in[n]$ then shows that ${\mathcal{C}}_{C}(\Omega_{\sigma},{\mathbf{u}})$ is integrally equivalent to $m{\mathcal{Q}}_{\sigma}$. Theorem 5 in [12] expresses the number of lattice points of ${\mathcal{C}}_{C}(\Omega_{\sigma},{\mathbf{u}})$ as a sum over all linear extensions of $\Omega_{\sigma}$. Such a linear extension is uniquely determined by the positions, say $1\leq j_{1}<j_{2}<\cdots<j_{n}\leq 2n$, of $\beta_{1},\beta_{2},\dots,\beta_{n}$, which must satisfy $j_{s_{i}}\geq 2s_{i}$ for every $i\in[k]$. Thus, setting $\kappa_{i}=j_{i}-j_{i-1}-1$ for $i\in[n]$ (where $j_{0}=0$), the argument of [12, p. 617] shows that in our situation, [12, Theorem 5] yields the formula
@@
where $K_{n}$ consists of all tuples $(\kappa_{1},\kappa_{2},\dots,\kappa_{n})\in{\mathbb{N}}^{n}$ such that $\kappa_{1}+\kappa_{2}+\cdots+\kappa_{n}=n$ and $\kappa_{1}+\kappa_{2}+\cdots+\kappa_{s_{i}}\geq s_{i}$ for every $i\in[k]$.
@@
for every $i\in[k]$ and ${\mathcal{Z}}(P_{{\rm rev}(\sigma)},m+1)$ is equal to the sum of the number of $m$-multichains in ${\mathbb{N}}^{n}$ with maximum element $(\lambda_{1},\lambda_{2},\dots,\lambda_{n})$ over all such tuples $(\lambda_{0},\lambda_{1},\dots,\lambda_{n})\in{\mathbb{N}}^{n+1}$. We may set $\kappa_{1}=\lambda_{0}+\lambda_{1}$ and $\kappa_{i}=\lambda_{i}$ for all $1<i\leq n$ and sum instead the number of $m$-multichains in ${\mathbb{N}}^{n}$ with maximum element $(\lambda_{1},\kappa_{2},\dots,\kappa_{n})$, where $0\leq\lambda_{1}\leq\kappa_{1}$, over all tuples $(\kappa_{1},\kappa_{2},\dots,\kappa_{n})\in K_{n}$. By Lemma 3.1 and basic enumerative combinatorics, the number of these multichains is equal to the right-hand side of Equation (11) and the proof follows. ∎
@@
(12) $\sum_{m\geq 0}{\rm Ehr}({\mathcal{Q}},m)t^{m}=\frac{h^{\ast}({\mathcal{Q}},t)}{(1-x)^{n+1}}.$
@@
(13) $h^{\ast}({\mathcal{Q}}_{\sigma},t)=\sum_{w}t^{{\rm asc}(w)},$
where $w$ ranges over all words $w\in[n]^{n}$ which have at least $s_{i}$ entries less than or equal to $s_{i}$ for every $i\in[k]$. Equivalently, the unique nondecreasing rearrangement $(u_{1},u_{2},\dots,u_{n})$ of $w$ should satisfy $u_{s_{i}}\leq s_{i}$ for every $i\in[k]$.
The proof employs the basics of the theory of EL-shellability of posets [4] [13, Lecture 3], which we briefly recall. Let $(P,\preceq)$ be a finite graded poset which has a minimum element $\hat{0}$ and a maximum element $\hat{1}$ and let ${\mathcal{E}}(P)$ be the set of covering relations of $P$. An edge labeling of $P$ (with integer labels) is any map $\lambda:{\mathcal{E}}(P)\rightarrow{\mathbb{Z}}$. Let $C$ be an unrefinable chain $p_{0}\prec p_{1}\prec\cdots\prec p_{m}$ of elements of $P$, so that $(p_{i-1},p_{i})\in{\mathcal{E}}(P)$ for every $i\in[m]$. We define the label of $C$ with respect to $\lambda$ as the sequence
@@
(see, for instance, [2, Equation (8)]) and ${\rm Ehr}({\mathcal{Q}}_{\sigma},t)={\mathcal{Z}}(P_{\tau},t+1)$ by Theorem 1.3, we have $h^{\ast}({\mathcal{Q}}_{\sigma},t)=h(\Delta(P_{\tau}),t)$. Thus, it suffices to show that $h(\Delta(P_{\tau}),t)$ is equal to the sum on the right-hand side of Equation (13).
@@
Proposition 3.2 specializes to a well known combinatorial interpretation of the $h^{\ast}$-polynomial (equivalently, of the $h$-polynomial of the $n$-fold edgewise subdivision) of the standard $n$-dimensional simplex when $\sigma=(n)$ has one part; see, for instance, [8, Corollary 8] for a generalization and refinement. A generalization of Proposition 3.2, valid for the polytopes in ${\mathbb{R}}^{n}$ defined by the equations

### v3

@@ 4. Ehrhart and Zeta polynomials
This section offers two proofs of Theorem 1.2 and deduces a combinatorial interpretation of the $h^{\ast}$-polynomial of the composition polytope ${\mathcal{Q}}_{\sigma}$. As in the previous sections, $\sigma=(r_{1},r_{2},\dots,r_{k})$ is a composition of $n$ and $s_{i}=r_{1}+r_{2}+\cdots+r_{i}$ for $0\leq i\leq k$, where $s_{0}:=0$.
@@
For any $m\in{\mathbb{N}}$, ${\rm Ehr}({\mathcal{Q}}_{\sigma},m)$ is equal to the number of points $(a_{1},a_{2},\dots,a_{n})\in{\mathbb{N}}^{n}$ such that
@@
Part of our second proof of Theorem 1.3 follows the proof of [16, Equation (33)], which is a formula for the number of lattice points of a Pitman–Stanley polytope. To avoid making too long of a digression, we assume familiarity with Sections 3 and 4 of [16] and only sketch this part of the proof.
@@
Let us consider the poset $\Omega_{\sigma}$ which consists of two disjoint chains $\alpha_{1}\prec\alpha_{2}\prec\cdots\prec\alpha_{n}$ and $\beta_{1}\prec\beta_{2}\prec\cdots\prec\beta_{n}$, together with the additional cover relations $\alpha_{s_{i}}\prec\beta_{s_{i}}$ for every $i\in[k]$. We let $C=\{\beta_{1},\beta_{2},\dots,\beta_{n}\}$ and ${\mathbf{u}}=(u_{1},u_{2},\dots,u_{n})$, where $u_{j}=jm$ for every $j\in[n]$ and where $m\in{\mathbb{N}}$ is fixed. In this situation, Pitman and Stanley [16, Section 3] have defined the order cone section ${\mathcal{C}}_{C}(\Omega_{\sigma},{\mathbf{u}})$ as the polytope of all order preserving maps $f:\Omega_{\sigma}\rightarrow{\mathbb{R}}_{\geq 0}$ such that $f(\beta_{j})=u_{j}$ for every $j\in[n]$. Setting $f_{j}=f(\alpha_{j})$ for every $j\in[n]$ and $f_{0}=0$ we have
@@
for every $i\in[k]$. The transformation $x_{j}=f_{j}-f_{j-1}$ for $j\in[n]$ then shows that ${\mathcal{C}}_{C}(\Omega_{\sigma},{\mathbf{u}})$ is integrally equivalent to $m{\mathcal{Q}}_{\sigma}$. Theorem 5 in [16] expresses the number of lattice points of ${\mathcal{C}}_{C}(\Omega_{\sigma},{\mathbf{u}})$ as a sum over all linear extensions of $\Omega_{\sigma}$. Such a linear extension is uniquely determined by the positions, say $1\leq j_{1}<j_{2}<\cdots<j_{n}\leq 2n$, of $\beta_{1},\beta_{2},\dots,\beta_{n}$, which must satisfy $j_{s_{i}}\geq 2s_{i}$ for every $i\in[k]$. Thus, setting $\kappa_{i}=j_{i}-j_{i-1}-1$ for $i\in[n]$ (where $j_{0}=0$), the argument of [16, p. 617] shows that in our situation, [16, Theorem 5] yields the formula
@@
where $K_{\sigma}$ consists of all tuples $(\kappa_{1},\kappa_{2},\dots,\kappa_{n})\in{\mathbb{N}}^{n}$ such that $\kappa_{1}+\kappa_{2}+\cdots+\kappa_{n}=n$ and $\kappa_{1}+\kappa_{2}+\cdots+\kappa_{s_{i}}\geq s_{i}$ for every $i\in[k]$.
@@
for every $i\in[k]$ and ${\mathcal{Z}}(P_{{\rm rev}(\sigma)},m+1)$ is equal to the sum of the number of $m$-multichains in ${\mathbb{N}}^{n}$ with maximum element $(\lambda_{1},\lambda_{2},\dots,\lambda_{n})$ over all such tuples $(\lambda_{0},\lambda_{1},\dots,\lambda_{n})\in{\mathbb{N}}^{n+1}$. We may set $\kappa_{1}=\lambda_{0}+\lambda_{1}$ and $\kappa_{i}=\lambda_{i}$ for all $1<i\leq n$ and sum instead the number of $m$-multichains in ${\mathbb{N}}^{n}$ with maximum element $(\lambda_{1},\kappa_{2},\dots,\kappa_{n})$, where $0\leq\lambda_{1}\leq\kappa_{1}$, over all tuples $(\kappa_{1},\kappa_{2},\dots,\kappa_{n})\in K_{\sigma}$. By Lemma 4.1 and basic enumerative combinatorics, the number of these multichains is equal to the right-hand side of Equation (13) and the proof follows. ∎
@@
(14) $\sum_{m\geq 0}{\rm Ehr}({\mathcal{Q}},m)t^{m}=\frac{h^{\ast}({\mathcal{Q}},t)}{(1-t)^{n+1}}.$
@@
(15) $h^{\ast}({\mathcal{Q}}_{\sigma},t)=\sum_{w\in{\mathcal{W}}_{\sigma}}t^{{\rm asc}(w)},$
where ${\mathcal{W}}_{\sigma}$ consists of all words $w\in[n]^{n}$ which have at least $s_{i}$ entries less than or equal to $s_{i}$ for every $i\in[k]$. Equivalently, the unique nondecreasing rearrangement $(u_{1},u_{2},\dots,u_{n})$ of $w$ should satisfy $u_{s_{i}}\leq s_{i}$ for every $i\in[k]$.
The proof employs the basics of the theory of EL-shellability of posets [5] [17, Lecture 3], which we briefly recall. Let $(P,\preceq)$ be a finite graded poset which has a minimum element $\hat{0}$ and a maximum element $\hat{1}$ and let ${\mathcal{E}}(P)$ be the set of covering relations of $P$. An edge labeling of $P$ (with integer labels) is any map $\lambda:{\mathcal{E}}(P)\rightarrow{\mathbb{Z}}$. Let $C$ be an unrefinable chain $p_{0}\prec p_{1}\prec\cdots\prec p_{m}$ of elements of $P$, so that $(p_{i-1},p_{i})\in{\mathcal{E}}(P)$ for every $i\in[m]$. We define the label of $C$ with respect to $\lambda$ as the sequence
@@
(see, for instance, [3, Equation (8)]) and ${\rm Ehr}({\mathcal{Q}}_{\sigma},t)={\mathcal{Z}}(P_{\tau},t+1)$ by Theorem 1.3, we have $h^{\ast}({\mathcal{Q}}_{\sigma},t)=h(\Delta(P_{\tau}),t)$. Thus, it suffices to show that $h(\Delta(P_{\tau}),t)$ is equal to the sum on the right-hand side of Equation (15).
@@
Proposition 4.2 specializes to a well known combinatorial interpretation of the $h^{\ast}$-polynomial (equivalently, of the $h$-polynomial of the $n$-fold edgewise subdivision) of the standard $n$-dimensional simplex when $\sigma=(n)$ has one part; see, for instance, [12, Corollary 8] for a generalization and refinement. A generalization of Proposition 4.2, valid for the polytopes in ${\mathbb{R}}^{n}$ defined by the equations
@@
Acknowledgements. The author wishes to thank Frederic Chapoton for his crucial input to the proof of Conjecture 1.1, described in Remark 3.2, and an anonymous reviewer for carefully reading the manuscript.